Circular No. 6857

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 AND CI CAMELOPARDALIS F. E. Marshall and T. E. Strohmayer, Goddard Space Flight Center;

and W. H. G. Lewin, Massachusetts Institute of Technology, on behalf of his collaborators, report a new position for XTE J0421+560 (IAUC 6855) based on multiple scans across the source with the Proportional Counter Array of RXTE on Apr. 1.3 UT. The best-fit position is α = 4^{h}19^{m}.6, δ = +56^{o}00′ (equinox 2000.0), with an estimated error radius (90-percent confidence) of 1′. This is 12′ from the previously reported position.

R. M. Hjellming and A. J. Mioduszewski, National Radio Astronomy Observatory, report: “On Apr. 2.63 UT, a 19-mJy radio source was observed with the Very Large Array at 1.4 GHz at α = 4^{h}19^{m}42^{s}.05 ± 0^{s}.03, δ = +55^{o}59′58′′.6 ± 0′′.5 (equinox 2000.0). This is inside the RXTE PCA position reported by Marshall et al. (see above). This source was detected on Apr. 1.9 UT at a level of ∼ 12 mJy at 1.4 GHz, but with poor flux determination due to bandwidth smearing and beam degradation, since it was 11′ from the field center. This source was not present in an NVSS survey image from 1993 Nov. 3 (upper limit of 2 mJy at the same frequency). Although this source could be the radio counterpart of XTR J0421+560 if short timescale variability can be established, R. M. Wagner (see below) has informed us that its position is coincident with that of the symbiotic star, CI Cam (= MWC 84).”

R. M. Wagner, Ohio State University; and S. G. Starrfield, Arizona State University, report: “An optical spectrum (range 355–715 nm; resolution 1.4 nm) of CI Cam was obtained on Apr. 3.145 UT with the Perkins 1.8m telescope and CCD spectrograph. The spectrum exhibits strong Balmer and He I emission lines, as well as weaker emission features arising from Fe II. For Hα emission, we find an instrument-corrected FWHM of 610 km/s, a FWZI of 2900 km/s, and an equivalent width of 32 nm. The Hα to Hβ intensity ratio is 6.4. Our spectrum is similar to that obtained in Jan. 1984 by Downes (1984, PASP 96, 807) and described previously by Merrill (1933, Ap.J. 77, 44). However, our spectrum also exhibits He II emission lines at 468.3 and 541.1 nm, with equivalent widths of 5.4 and 0.4 nm, respectively. He II emission lines are not present in the Downes spectrum, nor were they detected by Merrill. The presence of He II at this time suggests that CI Cam could be the optical counterpart of XTE J0421+560, in analogy with the optical spectra of other x-ray sources. Its appearance could also result from an orbital effect and be unrelated to the new x-ray source. Continued optical spectroscopy to monitor the evolution of the spectrum and He II emission is urged.”

1998 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green